OpenStudy (anonymous):

Can anyone help me with this problem? Simplify the following 2a^4b^9c^15/abc 2a^4b^9c^15 divided by abc Thanks!

geerky42 (geerky42):

\[\Large\dfrac{2a^4b^9c^{15}}{abc} = 2a^{4-1}b^{9-1}c^{15-1} = 2a^3b^8c^{14}\]

OpenStudy (anonymous):

Thanks!